# Identification of Craving Maps among Marijuana Users via the Analysis of Functional Brain Networks with High-Order Attention Graph Neural Networks

### v4

@@ 3.4 Ablation study
    We compared the performance solely using $\mathcal{L}_{c}$. from the experimental results. Our findings demonstrate that HOGANN, which integrates two sub-models, achieved an AUC of 78.8% and 66.1% in two datasets, respectively. Furthermore, using only $\mathcal{L}_{MixGAT}$ for the high-order attention effect showed no difference from $\mathcal{L}_{c}$, and its accuracy, precision, and recall were lower than when using only $\mathcal{L}_{c}$ in marijuana-323. It is evident that $\mathcal{L}_{c}$ can still preserve the optimization effects of high- order attention within a single loss function.
@@
    We believe that combining and optimizing loss functions may be better than using a single loss function. However, our experiment using a weight of $\alpha=0.3$ did not improve the results. Rather, it simply added the the two loss functions $\mathcal{L}_{c}+\mathcal{L}_{MixGAT}$. This indicates that HOGANN indeed improves optimization by combining the two losses, resulting in a classification AUC of 84.0%, an accuracy of 78.4%, a precision of 78.3%, and a recall of 78.4% in marijuana-323. Similar improvements in performance were seen in HCP.
@@
    In Figure 8, we increase the number of graph attention multi-heads to capture long-range dependencies and local neighborhood graph structure. From the results, it can be observed that HOGANN in marijuana-323 and HCP substantially improved classification performance. This indicates that increasing the number of attention heads can enable a broader learning of node embeddings for the high-order connectivity of each vertex.
@@
    We analyzed the HCP dataset characterized by marijuana dependence (Mj) and those who have previously used marijuana, conducting predictions across different age groups. The experimental results indicate that for the group with marijuana dependence, our HOGANN predicts the results more accurately in all age groups compared to the group that has only used marijuana. This indicates that individuals with marijuana dependence exhibit stronger brain region activity in the BOLD fMRI time series.
@@
Condition Age Group AUC (%) Accuracy (%)
Marijuana Dependence 22-25 88.5 92.4
25-30 89.5 86.4
30-37 98.8 87.9
Ever Used Marijuana 22-25 72.1 81.1
25-30 72.2 63.1
30-37 70.9 65.9
Figure 4: Comparing different models based on their classification performance on the marijuana-323 dataset with varying multi-segment time window sizes.
Figure 5: The average of predicted functional weighted connectivity (WC) and degree centrality (DC) values ranking between ROIs in marijuana-323. Panel (A) illustrates the comparison between the HC and LM groups, while Panel (B) displays the corresponding network of the HC and LM group, highlighting the regions with the highest DC values.
Figure 6: A comparison of machine learning methods with our HOGANN in identifying the optimal community clustering. In Panel (A), we compare community detection in weighted functional connectivity matrices using machine learning and our HOGANN, where we map the most significant four communities of the LM group to their corresponding brain regions of the connectivity network. Based on Panel (B) for the LM group, we depict four communities corresponding to the colored boxes, illustrating the connectivity of the most active brain regions in Panel (C).
Figure 7: Evaluating AUC performance using five-fold cross-validation on the marijuana-323 dataset for multigraph classification.
Figure 8: Panel (A) demonstrates the performance changes with multiple attention heads, while Panel (B) illustrates the classification performance changes when the neighborhood K in of the HOGANN is varied.
Table 3: Loss comparison
Marijuana-323 HCP
Strategy AUC (%) Acc. (%) Prec. (%) Rec. (%) AUC (%) Acc. (%) Prec. (%) Rec. (%)
$\mathcal{L}_{c}$ 78.8 76.9 76.7 76.9 66.1 63.9 63.9 63.9
$\mathcal{L}_{MG}$ 78.9 73.9 73.7 73.9 62.4 68.0 68.3 68.0
$\alpha\mathcal{L}_{MG}+(1-\alpha)\mathcal{L}_{c}$ 78.9 75.4 75.4 75.4 59.2 62.5 62.5 62.5
$\mathcal{L}_{c}+\mathcal{L}_{MG}$ 84.0 78.5 78.3 78.5 70.0 69.1 69.4 69.1
Table 4: Important Brain Connections in Marijuana-323
ROI index Connection Description
(1, 89) “L anterior insula” and “R crus cerebellum - visiospatial”.
(4, 89) “R anterior insula” and “R crus cerebellum - visiospatial”.
(6, 89) “R crus cerebellum - ant_sal” and “R crus cerebellum - visiospatial”.
(7, 89) “L superior temporal/auditory” and “R crus cerebellum - visiospatial”.
(10, 89) “R striatum/thalamus” and “R crus cerebellum - visiospatial”.
(12, 89) “L inferior frontal gyrus” and “R crus cerebellum - visiospatial”.
(16, 27) “L lateral angular gyrus” and “L mid-temporal cortex”.
(22, 89) “L parahippocampal gyrus” and “R crus cerebellum - visiospatial”.
(24, 89) “L mid occipital cortex” and “R crus cerebellum - visiospatial”.
(26, 89) “L inferior frontal gyrus” and “R crus cerebellum - visiospatial”.
(28, 89) “L mid-posterior temporal cortex” and “R crus cerebellum - visiospatial”.
(29, 89) “L medial angular gyrus” and “R crus cerebellum - visiospatial”.
(30, 89) “R inferior frontal gyrus” and “R crus cerebellum - visiospatial”.
(31, 89) “R mid-temporal cortex” and “R crus cerebellum - visiospatial”.
(52, 89) “bilateral medial posterior precuneus” and “R crus cerebellum - visiospatial”.
(55, 89) “bilateral calcarine cortex” and “R crus cerebellum - visiospatial”.
(56, 89) “L LGN” and “R crus cerebellum - visiospatial”.
(63, 64) “L pre/post-central gyri” and “R pre/post-central gyri”.
(67, 89) “cerebellar vermis” and “R crus cerebellum - visiospatial”.
(81, 89) “L precentral/fronto-opercular region” and “R crus cerebellum - visiospatial”.
Figure 9: The top three LM groups activate brain regions with mean DC values across four communities in the Marijuana-323.
Figure 10: HOGANN predicted craving location in the brain (green circle) using carving and gray regions of craving maps, Achieving high accuracy in determining overlaps related to chronic marijuana use in marijuana-323.
@@ 3.5 Community clustering
This study explores the community structure within brain networks, emphasizing the significance of clustering for each region based on reconstructed brain networks from Eq. (14). Our analysis examines the intricate connections between LM users and HC in the brain. This examination is facilitated by integrating the functional connectivity of each instance with the trainable weighted matrix derived from our HOGANN. Furthermore, we can rank the modularity and degree centrality (DC) by utilizing the average trainable weighted connectivity matrix to identify distinct clusters and community regions and compare with ML and our proposed model as demonstrated in Fig. 5. Further, we can find that the LM groups in Fig. 5 (A) have more significant community clustering, and we find out the significant region of the top five DC values in Fig. 5 (B).
We conducted a comparative analysis between traditional linear machine learning methods and our proposed HOGANN. This comparison, illustrated in Fig. 6 (linear methods) and B (HOGANN), highlights the superiority of the HOGANN in achieving enhanced modularity and clustering effects, particularly in community detection, as substantiated by the results.
@@
As depicted in Fig. 6 (C), we encircled squares based on the color of each brain cluster region, representing their high-density degree of centrality. By identifying critical brain regions associated with four communities and sorting the importance scores according to the average degree centrality values, we highlighted them. This allowed us to pinpoint essential brain regions linked to four distinct communities. In Table 4, we have followed the indices corresponding to 90 ROIs from Stanford Atlas definition  [36] and calculate DC scores that can be found within the four communities, as shown in Fig. 9. It reveals the importance of the rankings of these values, offering insights into the structural and functional connectivity across these brain regions. In our section 3.7 on brain analysis, we will analyze and interpret the subnetworks corresponding to the four communities through Yeo’s seven subnetworks.
@@ 3.6 Craving maps characterization
Craving maps can aid in deciphering brain network alterations in long-term marijuana users. We utilized the Neurosynth package (https://neurosynth.org/), aligning it with the ”craving” keyword to identify consistently active brain regions across studies strongly associated with the term “craving” [41]. Leveraging the design of the cited study, an automated meta-analysis of 80 publications identified regions activated by ”craving,” and Stanford ROIs can be mapped to the uniformity map  [33]. Specifically, the uniformity map measures low-activity voxels by applying a threshold of 25% during processing and identifies the activity in each brain region. The identified ROIs on the resulting map can be marked in red as threshold activity areas, as shown in Fig. 10. These ROIs represent areas substantially associated with craving, offering a comprehensive view of the brain’s response to LM use.
Next, we utilized the HOGANN to predict craving regions based on degree centrality (depicted by the green circle). We use uniformity maps to compare these predicted regions to the Neurosynth meta-analytic database. Finally, we invert the y-axis on the brain volumes within the Stanford ROIs, as indicated by the gray circle. The outcomes shown in Fig. 10 demonstrate the concordance of our HOGANN in predicting green cravings that are closely aligned with the ROIs across the axes at y = -60, y = 50, y = 32, y = 44 and y = 58. Our results showcase a successful analysis of cognitive-behavioral regions in LM users by effectively predicting specific activity regions. The prediction results through our model improve the prediction consistency and concordance of predictions between the network region and the craving maps from the meta-analysis of existing studies.
@@ 3.7 Alterations in the intrinsic functional connectivity
Yeo et al. [42] identified seven intrinsic functional brain networks of the cerebral cortex using the cluster approach based on resting state fMRI data from 1000 participants [42]. We utilized Yeo’s atlas of seven brain networks and analyzed the signal changes among these networks for SUD patients and normal controls. The seven networks include: 1. Visual attention 2. Somatosensory 3. Dorsal attention 4. Ventral attention 5. Limbic 6. Frontoparietal 7. Default mode network. In Fig. 11, we plotted a uniformity map based on Yeo 7 resting-state subnetworks and marked regions predicted by HOGANN (in blue). In the first row, the green regions correspond to overlapping subnetworks at the top of the atlas, which are areas of the Frontoparietal network associated with cognitive functions such as decision-making, problem-solving, planning, and executive control. In the second row, purple regions represent dorsal attention (attention and cognitive control). With Yeo’s seven subnetworks, we found the HOGANN could identify long-term marijuana users with substantial regions in an active transition of cognitive function and attention.
By comparing Yeo’s seven subnetworks to the four communities in Fig. 9, we can see that mean DC well explains the top three regions of each community. Communities 1 and 2 correspond to the dorsal attention network. Communities 4 and 5 correspond to the frontoparietal control network. Finally, Community 3 corresponds to regions that are either part of or closely related to the default mode network. According to our study’s observations, we align with previous relevant research that demonstrates LM users exhibit increased subcortical hyperconnectivity, leading to a widespread increase in functional connectivity across the whole brain connectome, particularly in the dorsal attention network and the default mode network [15, 43].

### v5

@@ 3.1 Ablation studies
    We evaluated model performance solely using loss function $\mathcal{L}_{c}$. Our experiments showed that HOGANN, which relied on a single model, achieved AUCs of 0.79 and 0.66 on the two datasets shown in Table 3. Notably, using only the $\mathcal{L}_{MixGAT}$ loss for high-order attention yielded similar performance to $\mathcal{L}{c}$ alone but with lower accuracy, precision, and recall in the marijuana-323 dataset. This suggests that the $\mathcal{L}_{c}$ loss effectively leverages the fused embedding effect of $\textbf{H}^{c}$. Although combining loss functions might seem beneficial, our experiment with a weighted sum loss ($\alpha=0.3$) of $\mathcal{L}{c}$ and $\mathcal{L}_{MixGAT}$ did not improve performance. Instead, HOGANN adds the two losses significantly improved results, achieving an AUC of 0.84, accuracy of 78.5%, precision of 78.3%, and recall of 78.5% in marijuana-323. Similar improvements in performance were seen in HCP.
@@
    In Fig. 3, we illustrate the Kullback-Leibler divergence (KLD) of two embeddings $\textbf{H}^{a}$ and $\textbf{H}^{\prime}$ from the HOGANN sub-models during the training process. Fig. 3 (left) depicts that HOGANN, when combined with MixHop, effectively stabilizes in KLD convergence of the two embeddings’ distributions during training process. Additionally, we also analyze Fig. 3 (right), which shows that the weighted losses exhibit more significant variation in the early stages of training. From our experiments, we observe that the weighted loss disperses the joint embedding of the two sub-models. The weighting also disperses the convergence of the individual sub-models.
@@
    In Fig. 4, we increase the number of graph attention layers with multi-head to capture long-range dependencies and local neighborhood graph structure, as illustrated in Fig. 4 (A), and (B). From the results, it can be observed that HOGANN in marijuana-323 and HCP substantially improved classification performance. This indicates that increasing the number of attention heads can enable a broader learning of node embeddings for the high-order connectivity of each vertex.
@@
    To evaluate classification based on varying lengths of fMRI time series, our experiments illustrate the impact of segmenting full-time series and subsequences into multigraphs with different window sizes $T^{\prime}$ in marijuana-323, as illustrated in Fig. 4 (C). Our training strategy employs four-time windows of length $T^{\prime}\in$0, 100, 300, 600} for training and testing. As depicted in Fig. 4 (C), it is evident that the AUC, accuracy, F1, and precision exhibit substantial reductions as the number of segmented fMRI signals increases (segment $>$ 3) in the marijuana-323. However, GAT demonstrates a more stable performance when segmenting two or three times compared to GCN. HOGANN showcases increasingly robust classification performance in our experiments as the number of segments increases. Specifically, under the optimal condition of segmenting fMRI signals twice, HOGANN achieves an AUC of 0.83, accuracy of 79.2%, F1 of 79.26%, and precision of 79.3%.
  5. 5.
    We analyzed the HCP dataset characterized by marijuana dependence (Mj) and those who have previously used marijuana, conducting predictions across different age groups. As shown in Table 4, the experimental results indicate that for the group with marijuana dependence, our HOGANN predicts the results more accurately in all age groups compared to the group that has only used marijuana. This indicates stronger deviations in neural networks in individuals with marijuana dependence versus marijuana users as compared to healthy control participants.
@@
Condition Age Group AUC Accuracy (%)
Marijuana Dependence 22-25 0.89 92.4
25-30 0.90 86.4
30-37 0.99 87.9
Ever Used Marijuana 22-25 0.72 81.1
25-30 0.72 63.1
30-37 0.71 65.9
Figure 6: Meta-analysis comparing HOGANN predictions with Neurosynth craving maps in marijuana-323. We superimposed 90 Stanford ROIs onto Neurosynth uniformity maps to identify craving-associated brain areas (shown in gray). The 3D voxel and average signal overlap illustrates regions of significant activation, with an applied z-score threshold (depicted in red) highlighting areas of substantial brain activity across studies focused on the keyword ‘craving’. The green highlighted areas indicate regions where HOGANN highly accurately predicted degree centrality, showing significant overlap with Stanford ROIs associated with chronic marijuana use and craving-related regions.
@@
Figure 8: The average of predicted functional weighted connectivity (WC) and degree centrality (DC) values ranking between ROIs in marijuana-323. Panel (A) illustrates the comparison between the HC and LM groups, while Panel (B) displays the corresponding network of the HC and LM groups, highlighting the regions with the highest DC values significant differences at paired sample $t$-test ($\star$” denotes $p<$ 0.05) in Marijuana-323.
@@ 3.2 Community clustering
This study explores the community structure within brain networks, emphasizing the significance of clustering for each region based on AWFC matrix and removing small connectivity edges with thresholds lower than 2%. Our analysis is to identify the intricate connections in the LM group brain network and find community partitions. This examination is facilitated by integrating the functional connectivity of each instance with the trainable weighted matrix derived from our HOGANN prediction. Furthermore, we can rank the modularity [42, 43] and DC by utilizing the AWFC matrix to identify distinct clusters and community regions and compare with ML and our proposed model as demonstrated in Fig. 5. This comparison, illustrated in Fig. 5 (A), demonstrates the derivation of the AWFC matrix using linear methods. The matrix is obtained by averaging the weight coefficients from four linear machine learning models (LR+L1, LR+L2, SVM+L1, and SVM+L2) and then multiplying these averages coefficient weight across all subjects in out-of-sample testing to find community clusters. In contrast, our proposed HOGANN-predicted AWFC matrix, as shown in Fig. 5 (B), highlights the superiority of the HOGANN in achieving enhanced modularity and has more significant community clustering, and we find out the significant region, particularly in community detection, as substantiated by the results.
@@
As depicted in Fig. 5 (C), we encircled squares based on the color of each brain cluster region, representing their high-density degree of centrality. We highlighted critical brain regions associated with four communities by identifying them and sorting the importance scores according to the average degree centrality values. This allowed us to pinpoint essential brain regions linked to four distinct communities. In Marijuana-323 dataset, we have followed the indices corresponding to 90 ROIs from Stanford atlas definition [39] and calculate DC scores that can be found within the four communities, as shown in Fig. 5 (D). It reveals the importance of the rankings of these values, offering insights into the spatial relationships in functional connectivity across these brain regions. In our section 3.4 on brain analysis, we will further analyze and interpret the regions of Yeo’s seven subnetworks corresponding to the four communities in Marijuana-323 (90 ROIs) and HCP (22 ROIs).
Figure 9: Visualization of mean DC values across seven distinct networks in Marijuana-323, based on diversity analysis of LM groups.
Table 5: Comparison of models predicting top-N Dice scores for keywords of Craving, Addiction, and Substance in target regions
Meta-Analysis Craving (num. studies=80) Addiction (num. studies=135) Substance (num. studies=124)
Top-15 Dice Top-20 Dice Top-30 Dice Top-15 Dice Top-20 Dice Top-30 Dice Top-15 Dice Top-20 Dice Top-30 Dice
LR + L1 0.27 0.29 0.36 0.18 0.26 0.33 0.31 0.27 0.34
LR + L2 0.21 0.24 0.32 0.18 0.26 0.33 0.30 0.26 0.38
SVM + L1 0.21 0.30 0.37 0.19 0.28 0.34 0.31 0.27 0.34
SVM + L2 0.21 0.24 0.33 0.13 0.22 0.30 0.26 0.22 0.30
XGboost 0.27 0.28 0.40 0.19 0.32 0.38 0.32 0.28 0.30
HOGANN 0.39 0.44 0.48 0.22 0.35 0.40 0.33 0.34 0.47
Table 6: Brain regions categorized by affiliated Functional subnetwork.
ROIs Frontoparietal Network (FPN) Default Mode Network (DMN) Dorsal Attention Network (DAN)
Stanford Top-5 region R middle frontal/dlPFC, L middle frontal/dlPFC bilateral ACC L inferior parietal/angular gyrus
Top-10 region R middle frontal/dlPFC, L middle frontal/dlPFC bilateral ACC, bilateral medial posterior precuneus, bilateral mPFC, bilateral posterior cingulate L inferior parietal/angular gyrus, R inferior parietal/angular gyrus
Top-15 region R middle frontal/dlPFC, L middle frontal/dlPFC, R middle frontal gyrus, L inferior frontal gyrus bilateral ACC, bilateral medial posterior precuneus, bilateral mPFC, bilateral posterior cingulate L inferior parietal/angular gyrus, R inferior parietal/angular gyrus
Glasser Top-5 region Orbital_and_Polar_Frontal, Inferior_Frontal Medial_Temporal, Lateral_Temporal, Posterior_Cingulate Primary_Visual, Dorsal_Stream_Visual, MT Complex and Neighboring Visual_Areas
Top-10 region Orbital_and_Polar_Frontal, Inferior_Frontal, Dorsolateral_Prefrontal Medial_Temporal, Lateral_Temporal, Posterior_Cingulate, Anterior_Cingulate, Medial_Prefrontal
Top-15 region Orbital_and_Polar_Frontal, Inferior_Frontal, Dorsolateral_Prefrontal Medial_Temporal, Lateral_Temporal, Posterior_Cingulate, Anterior_Cingulate, Medial_Prefrontal
@@ 3.3 Craving maps characterization
Craving maps can aid in deciphering brain network alterations in LM users. We utilized the Neurosynth package (https://neurosynth.org/), aligning it with the ”craving” keyword to identify consistently active brain regions across studies strongly associated with the term “craving” [44]. Leveraging the design of the cited study, an automated meta-analysis of 80 publications identified regions activated by ”craving,” and Stanford ROIs can be mapped to the uniformity map [34]. Specifically, the uniformity map measures low-activity voxels by applying a threshold of 25% during processing and identifies the activity in each brain region. The identified ROIs on the resulting map can be marked in red as threshold activity areas, as shown in Fig. 6. These ROIs represent areas substantially associated with craving, offering a comprehensive view of the brain’s response to LM use.
Next, we utilized the HOGANN to predict craving regions based on degree centrality (depicted by the green circle). We use uniformity maps to compare these predicted regions to the Neurosynth meta-analytic database. Finally, we invert the y-axis on the brain volumes within the Stanford ROIs, as indicated by the gray circle. The outcomes shown in Fig. 6 demonstrate the concordance of our HOGANN in predicting green cravings that are closely aligned with the ROIs across the axes at y = -60, y = -50 and y = 32. Our results showcase a successful analysis of cognitive-behavioral regions in LM users by effectively predicting specific activity regions. The prediction results through our model improve the prediction consistency and concordance of predictions between the network region and the craving maps from the meta-analysis of existing studies.
@@ 3.4 Alterations in the intrinsic functional connectivity
Yeo et al. [45] identified seven intrinsic functional brain networks of the cerebral cortex using the cluster approach based on fMRI data from 1000 participants [45]. We utilized Yeo’s atlas of seven brain networks and analyzed the signal changes among these networks for SUD patients and normal controls. The seven networks include: 1. Visual Network (VN) 2. Sensorimotor Network (SMN) 3. Dorsal Attention Network (DAN) 4. Ventral Attention Network (VAN) 5. Limbic Network (LN) 6. Frontoparietal Network (FPN) 7. Default Mode Network (DMN) 8. Subcortical System (SUB). In Fig. 7, we plotted a uniformity map based on Yeo 7 resting-state subnetworks and marked regions predicted by HOGANN (in blue). In the first row, the green regions correspond to overlapping subnetworks at the top of the atlas, which are areas of the DAN associated working memory task [46]. In the second row, red regions represent DMN (attention and cognitive control). With Yeo’s seven subnetworks, we found the HOGANN could identify LM users with substantial regions in an active transition of cognitive function and attention.
According to our study’s observations, we align with previous relevant research that demonstrates LM users exhibit increased subcortical hyperconnectivity, leading to a widespread increase in functional connectivity across the whole brain connectome, particularly in the DMN and partially in the FPN and DAN [16, 47]. Fig. 8 shows the 90 ROIs from the Stanford atlas definition  [39] and the calculated DC matrices between the HC and LM group. This finding was corroborated by our analysis of the distinct DC matrices for the HC and LM groups in Fig. 8 (A), which revealed a higher density of functional activity regions in the LM group compared to the HC group. Notably, the LM group exhibited increased density in the FPN, DMN, and DAN networks, as illustrated in Fig. 8 (B). Moreover, this observation aligns with and is significantly supported by previous research on functional connectivity changes in networks among 2,390 addiction patients (1,143 with behavioral addiction and 1,456 with substance use disorder SUD) and 2,610 HC. These earlier studies consistently identified changes primarily involving the DMN and FPN [48].
3.5 Craving region evaluation
To better evaluate the overlap of craving region between HOGANN prediction and activate brain maps, the maps were thresholded at p$<$0.01 with FDR correction, and we can define the top-$N$ Dice score as:
$Dice(N)=\frac{2|Prediction(N)\cap Target(N)|}{|Prediction(N)|+|Target(N)|}$ (23)
where $|Prediction(N)|$ and $|Target(N)|$ represent the number of top-N ROIs in the predicted activation map and the ground truth ROIs (e.g., Stanford 90 ROIs). The $|Prediction(N)|+|Target(N)|$ present the overlap between predicted Top-N region and target ROIs. Observing the results in Table 5, we analyzed keywords such as Craving, Addiction, and Substance under HOGANN. We compared their Top-N-Dice scores with other baseline methods from the meta-analysis. HOGANN showed the second-highest Dice score. For the keyword ‘Craving’ HOGANN’s Top-20 Dice score improved by 46.6%. For the keyword "Addiction," the Top-15 Dice score improved by 15.7%, and for "Substance," the Top-30 Dice score improved by 23.6%. The results indicate that HOGANN can identify better overlapping regions in the craving regions associated with LM use.
4 Discussion
This study is the first to investigate the spatial craving maps among the LM users and the prediction of brain functional connectivity alterations based on two fMRI datasets. Our findings reveal several critical insights into the brain’s functional organization associated with marijuana use: (1) HOGANN predicted the results more accurately for individuals with marijuana dependence versus marijuana users, demonstrating that these changes in brain function were linked to severity of use, (2) we used HOGANN to predict mean DC levels in different atlas regions (e.g., 90 and 22 ROIs) across the two datasets, focusing on the most frequently associated regions within Yeo’s seven networks. Notably, among the Top-N regions in these two atlases, the most frequently shared affiliated functional subnetworks were the FPN, DMN, and DAN, as shown in Table 6, (3) by comparing Yeo’s seven subnetworks to the four communities in Fig. 5 (D), community 1, 3 and 4 mainly involve the DMN, while community 2 and mainly involve the FPN. These results align with the changes in functional connectivity caused by nicotine use [49] and the significant impact on the DMN regions caused by marijuana administration [50], (4) compared to other ML methods, HOGANN demonstrated the highest Dice score for the keyword ’Craving’ among all keywords analyzed.
Several neurological studies have examined the treatment of marijuana addiction and its effects on three crucial brain networks including FPN, DMN, and DAN. These networks consistently show enhanced activity during drug-related cue processing but diminished activity in non-drug-related cognitive tasks [8]. FPN and DAN abnormalities appear to span across various stages of addiction, with FPN activity positively correlating with addiction severity, craving levels, and relapse risk, while DAN activity is associated with frequency of drug use. DMN abnormalities, on the other hand, are primarily associated with subjective craving. Importantly, the FPN is a crucial hub among brain networks [51] and therapeutic interventions have shown potential in upregulating FPN activity for addiction treatment [52, 53].
Moreover, we further examined how different age cohorts changed across time periods. Our analysis revealed that individuals aged 25-30 displayed greater mean DC in the DMN compared to the 22-25 age group. Notably, older participants demonstrated more pronounced alterations in both the DMN and VAN over time, as illustrated in Fig. 9. These findings suggest a positive relationship between advancing age and functional connectivity [54, 55, 56, 57].